# TokenMixer-Large: Scaling Up Large Ranking Models in Industrial Recommenders

### v1

@@ Abstract.
In recent years, the study of scaling laws for large recommendation models has gradually gained attention. Works such as Wukong, HiFormer, and DHEN have attempted to increase the complexity of interaction structures in ranking models and validate scaling laws between performance and parameters/FLOPs by stacking multiple layers. However, their experimental scale remains relatively limited. Our previous work introduced the TokenMixer architecture, an efficient variant of the standard Transformer where the self-attention mechanism is replaced by a simple reshape operation, and the feed-forward network is adapted to a pertoken FFN. The effectiveness of this architecture was demonstrated in the ranking stage by the model presented in the RankMixer paper. However, this foundational TokenMixer architecture itself has several design limitations. In this paper, we propose TokenMixer-Large, which systematically addresses these core issues: sub-optimal residual design, insufficient gradient updates in deep models, incomplete MoE sparsification, and limited exploration of scalability. By leveraging a mixing-and-reverting operation, inter-layer residuals, the auxiliary loss and a novel Sparse-Pertoken MoE architecture, TokenMixer-Large successfully scales its parameters to 7-billion and 15-billion on online traffic and offline experiments, respectively. Currently deployed in multiple scenarios at ByteDance, TokenMixer‑Large has achieved significant offline and online performance gains.
@@ 3.3.1. Mixing & Reverting
As shown, the number of tokens in the original input is $T$, while the number of tokens after mixing in the i-th layer is H. Unless each layer’s $H$ remains the same and $T$ is equal to $H$, residual connections cannot propagate smoothly across layers. To overcome this limitation, we design TokenMixer-Large, featuring a special two-layer TokenMixer structure: the first layer is responsible for mixing information among the original tokens, and the second layer specifically restores the dimension of the mixed tokens. This symmetric ”mixing–reverting” design ensures dimensional consistency of the input and output across modules, thereby establishing a continuous signal pathway from the initial input to the deep network. It enables stable residual connections at every layer and effectively prevents the loss of gradient information. Detailed comparison can be found in section 4.3
@@
(13) $[...,[\mathrm{x}_{t}^{(1)},\mathrm{x}_{t}^{(2)},\dots,\mathrm{x}_{t}^{(h)}],...]=\text{split}(\mathrm{\mathbf{H}^{next}})\in\mathbb{R}^{T\times H\times(D/H)}$
(14) $\mathrm{X_{t}^{revert}}=\text{concat}[\mathrm{x}_{t}^{(1)},\mathrm{x}_{t}^{(2)},\dots,\mathrm{x}_{t}^{(H)}]\in\mathbb{R}^{D}$
(15) $\mathrm{\mathbf{X}^{revert}}=\text{concat}[\mathrm{X_{1}^{revert}},\mathrm{X_{2}^{revert}},\dots,\mathrm{X_{h}^{revert}}]\in\mathbb{R}^{T\times D}$
@@ 3.4.1. First Enlarge, Then Sparse
In terms of iteration strategy, we adopt a “first enlarge, then sparse” approach: first scaling up the model to gain effectiveness, then splitting the SwiGLU networks and activating them sparsely to obtain efficiency benefits. The Sparse-Pertoken MoE (S-P MoE) can be formulated as follows:
@@
where $\{\mathrm{W}_{\text{up}}^{t},\mathrm{W}_{\text{gate}}^{t}\}\in\mathbb{R}^{D\times nD/E}$, $\{\mathrm{b}_{\text{up}}^{t},\mathrm{b}_{\text{gate}}^{t}\}\in\mathbb{R}^{nD/E}$, $\mathrm{W}_{\text{down}}^{t}\in\mathbb{R}^{nD/E\times D}$, $\mathrm{b}_{\text{down}}^{t}\in\mathbb{R}^{D}$. $E$ is the expert number of each token. $k$ is the top-k number for the router, in which the $\rm softmax(\cdot)$} and $\rm topk(\cdot)$} operations are used for the training and serving sparsity. Detailed design concept and experimental results can be found in Appendix A.1.
@@ 3.4.3. Gate Value Scaling
In practice, we observed that since $g(x)$ typically uses softmax to convert router logits into a probability distribution summing to 1, this sum-to-one constraint can lead to insufficient gradient updates for the pSwiGLUs. Therefore, we add a constant $\mathbf{\alpha}$ before $g(x)$ to allow more sufficient updates of the expert networks:

### v2

@@ Abstract.
While scaling laws for recommendation models have gained significant traction, existing architectures such as Wukong, HiFormer, and DHEN often struggle with sub-optimal designs and hardware under-utilization, limiting their practical scalability. Our previous TokenMixer architecture (introduced in the RankMixer paper) balanced effectiveness and efficiency by replacing self-attention with a lightweight token mixing operator; however, it encountered critical bottlenecks in deeper configurations, including sub-optimal residual paths, vanishing gradients, incomplete MoE sparsification, and constrained scalability. In this paper, we propose TokenMixer-Large, a systematically evolved architecture designed for extreme-scale recommendation. By introducing a mixing-and-reverting operation, inter-layer residuals, and an auxiliary loss, we ensure stable gradient propagation as model depth increases. Furthermore, we incorporate a Sparse Per-token MoE to enable efficient parameter expansion. TokenMixer-Large successfully scales to 7 billion parameters for online traffic and 15 billion in offline experiments. Currently deployed in multiple scenarios at ByteDance, TokenMixer-Large has achieved substantial performance gains: delivering a 1.66% increase in orders and a 2.98% boost in per-capita preview payment GMV for e-commerce; improving ADSS by 2.0% in advertising; and achieving 1.4% revenue growth in live streaming.
@@ 3.3.1. Mixing & Reverting
As shown, the number of tokens in the original input is $T$, while the number of tokens after mixing in the i-th layer is H. Unless each layer’s $H$ remains the same and $T$ is equal to $H$, residual connections cannot propagate smoothly across layers. To overcome this limitation, we design TokenMixer-Large, featuring a special two-layer TokenMixer structure: the first layer is responsible for mixing information among the original tokens, and the second layer specifically restores the dimension of the mixed tokens. This symmetric ”mixing–reverting” design ensures dimensional consistency of the input and output across modules, thereby establishing a continuous signal pathway from the initial input to the deep network. It enables stable residual connections at every layer and effectively prevents the loss of gradient information. Detailed comparisons with RankMixer(TokenMixer) can be found in section 4.3
@@
(13) $[...,[\mathrm{x^{\prime}}_{1}^{(h)},\mathrm{x^{\prime}}_{2}^{(h)},\dots,\mathrm{x^{\prime}}_{T}^{(h)}],...]=\text{split}(\mathrm{\mathbf{H}^{next}})\in\mathbb{R}^{T\times H\times(D/H)}$
(14) $\mathrm{X_{t}^{revert}}=\text{concat}[\mathrm{x^{\prime}}_{t}^{(1)},\mathrm{x^{\prime}}_{t}^{(2)},\dots,\mathrm{x^{\prime}}_{t}^{(H)}]\in\mathbb{R}^{D}$
(15) $\mathrm{\mathbf{X}^{revert}}=\text{concat}[\mathrm{X_{1}^{revert}},\mathrm{X_{2}^{revert}},\dots,\mathrm{X_{T}^{revert}}]\in\mathbb{R}^{T\times D}$
@@ 3.4.1. First Enlarge, Then Sparse
In terms of iteration strategy, we adopt a ”first enlarge, then sparse” approach: first scaling up the model to gain effectiveness, then splitting the pertoken SwiGLUs and activating them sparsely to obtain efficiency benefits. The Sparse-Pertoken MoE (S-P MoE) can be formulated as follows:
@@
where $\{\mathrm{W}_{\text{up}}^{t},\mathrm{W}_{\text{gate}}^{t}\}\in\mathbb{R}^{D\times nD/E}$, $\{\mathrm{b}_{\text{up}}^{t},\mathrm{b}_{\text{gate}}^{t}\}\in\mathbb{R}^{nD/E}$, $\mathrm{W}_{\text{down}}^{t}\in\mathbb{R}^{nD/E\times D}$, $\mathrm{b}_{\text{down}}^{t}\in\mathbb{R}^{D}$. $E$ is the expert number of each token. $k$ is the top-k number for the router $g(\cdot)$. And $g(\cdot)$ of the selected experts is constrained to a sum of 1 by $\rm softmax(\cdot)$}, representing that the sum of the probabilities of activated experts is 1. Detailed design concept and experimental results can be found in Appendix A.1.
@@ 3.4.3. Gate Value Scaling
In practice, we observed that since $g(\cdot)$ typically uses softmax to convert router logits into a probability distribution summing to 1, this sum-to-one constraint can lead to insufficient gradient updates for the SwiGLUs. Therefore, we add a constant $\mathbf{\alpha}$ before $g(\cdot)$ to allow more sufficient updates of the expert networks: